OpenStudy (anonymous):

You are choosing between to healthclubs. Club A offers members a fee of 70 plus a monthly fee of 25. Club B offers members fee of 45 plus a 45 monthly fee. 1.set up an equation for each club. 2. After how many months will total cost at each club will be the same. 3. What will be total cost for each club from part 2.

OpenStudy (anonymous):

What is the rate of change? the members fee or the monthly fee?

OpenStudy (anonymous):

Club B offers a fee of 45 plus a 30 monthly fee. Correction

OpenStudy (anonymous):

that's not what I asked.....

OpenStudy (anonymous):

The monthly fee

OpenStudy (anonymous):

Great so for Club A it will be. \[70+25x=y\] where x is the months and y is the total cost. Next for Club B we have \[45+45x=y\]

OpenStudy (anonymous):

That is part 1. Then you set the equations together for part 2 and solve for x. Remember x is the number of months.

OpenStudy (anonymous):

\[45+45x=70+25x\]

OpenStudy (anonymous):

Then after that you fort part 3 you simply plug back in the x=? you found in one of the two equations and solve for y

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

Dont know how to solve

OpenStudy (anonymous):

Do you know how to add divide and subtract?